OpenStudy (anonymous):

What is the radius of a communications satellite’s orbital path that is in a uniform circular orbit around Earth and has a period of exactly 24.0 hours (86,400 seconds). (Measurement is from the center of Earth.) A. 6.5 × 10^4 m B. 3.8 × 10^5 m C. 4.2 × 10^7 m D. 9.5 × 10^7 m

OpenStudy (anonymous):

 G = 6.67 × 10^-11 N.m2/kg2 ME = 5.98 × 10^24 kg 1 hr = 3,600 s

OpenStudy (souvik):

yeah...you are right...

OpenStudy (anonymous):

but i still dont know wat to do my teacher hasn't tached us this yet

OpenStudy (souvik):

use your calculator and calculate it...

OpenStudy (souvik):

in which class do you read?

OpenStudy (anonymous):

in physics AP and you

OpenStudy (souvik):

what is AP?

OpenStudy (anonymous):

Advanced placement :)

OpenStudy (souvik):

is it in secondary education or high school

OpenStudy (anonymous):

High school it's my last year in high school im a senior already :) what grgrade are you in

OpenStudy (souvik):

i just passed high school..

OpenStudy (anonymous):

really wow congrats how does it feel ?

OpenStudy (souvik):

missing those school days...school friends..:( college admission is in this month...

OpenStudy (anonymous):

really what college do you want to go to

OpenStudy (souvik):

i want to study physics honors....no engineering..

OpenStudy (souvik):

okay...close this question and post a new one...:)